OpenStudy (anonymous):

George had to borrow $36,000 for a new caf for 8months. The interest rate was 15% per year. How much interest did he gave for borrowing the $$ for 8months?

OpenStudy (anonymous):

i think it is compound interest

OpenStudy (anonymous):

A= P(1+(r/n))^(nt) P= $36,000 r=0.15 n= 1 t=8/12